OpenStudy (anonymous):

How would I prove the following: If a and b are real numbers such that a 1 + (k+1)a

OpenStudy (usukidoll):

isn't one of those questions Bernoullis?!

OpenStudy (anonymous):

im not sure what that is.. my professor gave us some exercises to work and im having trouble with those 2

OpenStudy (usukidoll):

one of them looks familiar... like http://en.wikipedia.org/wiki/Bernoulli%27s_inequality

OpenStudy (anonymous):

im not sure about that.. How would I prove it though?

OpenStudy (usukidoll):

idk...I'm in a proof writing class myself and my class is going crazy

OpenStudy (anonymous):

lol same here.. .-.

OpenStudy (anonymous):

if a < b, then by definition of <, b + addin(a) ∈ R+ let n = b + addin(a) n = n + 0 by additive identity n = n + b + addin(b), by additive inverse n = n + addin(addin(b)) + addin(b), by the lemma addin(addin(n)) = n n = b + addin(a) + addin(addin(b)) + addin(b), by substitution n = b + addin(b) + addin(a) + addin(addin(b)), by associative n = 0 + add(a) + addin(addin(b)), by additive inverse n = add(a) + addin(addin(b)), by identity property and since n∈ R+, and by definitio of <, add(b) < add(a) which means -b < -a man I suck -.-

OpenStudy (usukidoll):

Here's a great question. W T Fuu fuu invented pure math?

OpenStudy (anonymous):

humans XD

OpenStudy (anonymous):

I believe you can use binomial theorem for the second proof

OpenStudy (usukidoll):

oh no! X( torture

OpenStudy (usukidoll):

use induction XD

OpenStudy (anonymous):

yep, believe I got it

OpenStudy (anonymous):

base case k = 1 (1+a)^(1+1) >? 1 + (1+1)a (1+a)^2 >? 1 + 2a 1 + 2a + a^2 > 1+2a (check) suppose (1+a)^(k+1) > 1 + (k+1)a for for some integer k, consider (1+a)^(k+1+1) = (1+a) (1+a)^(k+1) by hypothesis: (1 + a)^(k+1) > 1 + (k+1)a since a > -1, then a+1 > 0, so we can multiply both sides by (1+a) without changing the sign (1+a) (1+a)^(k+1) > [1 + (k+1)a ] (1+a) expand the right side, > 1+ 2a + a^2+ ak + a^2 k > 1 + 2a + ak > 1 + (a + a + ak) = 1 + (1 + 1 + k) a check XD